The position of a body is given by x=at+4bt^3, where a and b are constants, x is position and t is time. Find velocity and acceleration at t=5seconds

Grade:11

2 Answers

Arun
5 years ago
Dear Dishita
 
x = at + 4bt^3
 
v = dx/ dt = a + 12 b t^2
 
a = dv / dt = d2x / dt2 = 24 bt
 
now velocity (t = 5 sec) = a + 300b
 
acceleration (t = 5 sec) = 120b
Khimraj
5 years ago
x = at + 4bt3
velocity is given by v = dx/dt = a + 12bt2
at t = 5,
v = a + 300b
and acceleration is given by
a = d2x/dt2 = dv/dt = 24bt
at t = 5
a = 120b.
Hope it clears.